OpenStudy (anonymous):

I am going to do a paper on how two magnets repel each other. I was thinking of having such an experiment where a pair of bar magnets are in a vertically placed tube and there is a certain distance between them because of the repulsion if they are correctly aligned. Could someone outline what are the factors taking place in the situation that definitively determine the distance between the bar magnets?

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):

the force of the floating magnet is the sum of the gravitational force (which points down) and the magnetic force ( which, for this alignment points up

OpenStudy (anonymous):

the opposite charges (negative and positive) never attract, causing the 2 magnets to repel from each other

OpenStudy (unklerhaukus):

the gravitational force on the floating magnet will be \(F_\text{gravitational}=mg\) and will act downwards. where \(g\) is acceleration dues to gravity and \(m\) is the mass of the mangent the force between pairs of magnetic poles will be \[F_\text{magnetic}=\frac{\mu}{4\pi}\frac{q_{m_1}q_{m_2}}{r^2}\] where \(\mu\) is about 1 in air and \(q_m\) is the strength of each pole, [if you use identical magnets \(q_{m_1}=q_{m_2}\)] \(r\) is the distance between the poles if the bar magnets are long enough the north poles in the middle will be the only poles you need to consider if you can measure the mass of the magnets and also measure \(r\) from experiment with this information you should be able to calculate the pole strength \(q_m\) of your magnets or if you know the pole strength already you should be able to predict the distance between the poles in this set up

OpenStudy (unklerhaukus):

also if you perform the same experiment in a tube of something other than air, then you will need to consider \(\mu\) the permeability of that substance